i am having trouble with functions. could you please help solve this question?
If f(x)=x^3-2x+1 then f(2/3)=

i figured -1/27 was the answer but i am not sure. thanks for your help.

Answer by checkley75(3666)

F(2/3)=(2/3)^3-2(2/3)+1
F(2/3)=(8/27)-4/3+1
F(2/3)=(8+9*-4+27)/27
F(2/3)=(8-36+27)/27
F(2/3)=-1/27
RIGHT ON!!!!!!!!!!!!!!! : )